what is the maximum number of 4-cm cubes that can fit into abox which measures 15 cm by 10 cm by 18 cm
Answer by Theo(13342)

You have:

height = 15 cm.
width = 10 cm.
depth = 18 cm.

Take 4 cm cubes and lay them in the bottom of the box.

You get 2 cubes wide by 4 cubes deep by 1 cube high.

You do this with 8 cubes.

Lay another layer.

that's another 8 cubes.

Lay another layer.

That's another 8 cubes.

All together you havre 24 cubes in the container.

That's 3 cubes * 2 cubes * 4 cubes = 24 cubes.

You can't put anymore in because 15 inch cube height / 4 = 3 plus a fraction, and 10 / 4 = 2 plus a fraction, and 18 / 4 = 4 plus a fraction, and fractions are not allowed.

Total square cm you have used is 12 * 8 * 16 = 1536.

Total square cm in the container is 15 * 10 * 18 = 2700.

The difference is empty space.